The length of a rectangle is
2

cm
longer than its width.
If the perimeter of the rectangle is
44

cm
, find its area.

Answer by addingup(3677)

2L + 2W = 44
L = W + 2
2(W + 2) + 2W = 44
2W + 4 + 2W = 44
4W = 40
W = 10 and L = 10 + 2 = 12
Area:
L x W = Area
12 x 10 = 120 cm^2 this is the area